28. For a gas in a closed container its pressure is found to be 400 N/m2 at 27°c. If the temperature were raised to 54°c, what would the pressure become? A) 800 N/m2 B) 436 N/m2 c) 200 N/m2 D) 427N /m2 The answer is B .Why ?
Answer #1

Given:

Pi = 400 N/m^2

Ti = 27.0 oC

= (27.0+273) K

= 300 K

Tf = 54.0 oC

= (54.0+273) K

= 327 K

use:

Pi/Ti = Pf/Tf

400 N/m^2 / 300.0 K = Pf / 327.0 K

Pf = 436 N/m^2

Answer: B